OpenStudy (anonymous):

The width of a rectangle is 8 more than the length. The perimeter of the rectangle is 56. What are the dimensions of the rectangle?

OpenStudy (anonymous):

A=L*W A=(L)(8+L) P=2W+2L P=2(8+L)+2L P=56 56=16+2L+2L 56=16+4L 40=4L L=10 A=(L)(8+L) A=(10)(8+10) A=180 The area is 180, the width is 18, and the length is 10. Does that make sense?

OpenStudy (littlebird):